OpenStudy (anonymous):

A single tree produces about 2.6 × 10^2 lb of oxygen each year. The Amazon rainforest has about 3.9 × 10^11 trees. About how many pounds of oxygen are produced in the rainforest each day ? 2.8 × 10 what is the exponent

OpenStudy (anonymous):

For you to find the number of pounds of oxygen produced you just have to divide 2.6x10^2 by the number of days per year and multiply it with the number of trees

OpenStudy (anonymous):

so would it be 676? @I_Need_Help_With_Home

OpenStudy (anonymous):

\[2.778x10^{11}\]

OpenStudy (anonymous):

soooo ten?

OpenStudy (anonymous):

or 11?

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

Its 11. and sorry I had to eat

OpenStudy (jordanloveangel):

1 tree = 2.6 * 10^2 lbs oxygen so 3.9* 10^11 times 1 tree= 3.9*10^11 * 2.6 * 10^2 lbs oxygen for the year yes it would be 3.9*2.6 * 10^(11+2 ) = this is produced in 1 year and the question asks for the day so you have the amount for 365 days and you need to know how much this is per day when you multiplied you got 10.14 * 10^13 = 1.014* 10^14 for the year divide by 365 to get for the day = 0.002778 * 10^14 2.778 * 10^ -3 *10^14 = 2.778 * 10^ (-3 + 14 )= 2.8 * 10^11 you would get the same result if you did for the day first and then multiplied by the number of trees. when you multiply numbers with the same base add the powers (exponents) when using scientific notation the base is 10 and the general form is a * 10^n where a is >0 and < 10 and n is an integer